# Search for the production of dark Higgs in the framework of Mono-Z portal at the FCC-ee simulated electron-positron collisions at $\sqrt{s}=240$ GeV

### v1

@@ Abstract
In the present work, we study the possible production of the dark Higgs boson ($h_{D}$) candidates, which originated from a simplified-model scenario based on the Mono-Z model, in association with a neutral gauge boson (Z). This study has been performed by studying events with dimuon plus missing transverse energy produced in the simulated electron-positron collisions at the foreseen Future Circular Collider in the Electron-Positron collision mode (FCC-ee), operating at 240 GeV center of mass energy and integrated luminosity of 10.8 ab-1. In case no new physics has been discovered, we set upper limits at a 95% confidence level on the mass of the dark Higgs.
arxiv: arXiv: V4
@@ 1 Introduction
If the Z does not interact with quarks, the HL-LHC and future hadron colliders will be unable to confirm its existence. In such a scenario, electron-positron colliders, like the proposed Future Circular Collider in its Electron-Positron collision mode (FCC-ee), will play a vital role. The FCC-ee is expected to begin operations at a center of mass energy of 240 GeV ($\sqrt{s}$), with plans for a future upgrade to 365 GeV FCC-ee1 ; FCC-ee2 ; FCC-ee3 .
The initial phase of the FCC-ee is designed to be a high-luminosity and high-precision electron-positron collider, functioning as a factory for electroweak processes, top quarks, and particularly, the Higgs boson FCC-ee4 . It provides controllable energy levels and minimizes QCD background noise, making it a crucial tool in this research area.
This analysis explores a light dark Higgs ($h_{D}$) which is produced at FCC-ee in addition to a light neutral gauge bosons (Z) with a mass $M_{Z^{\prime}}\leq 90$ GeV, based on the dark Higgs (DH) simplified model within the mono-Z framework R1 . We examine simulated electron-positron collisions at the FCC-ee with 240 GeV center of mass energy, focusing on dimuon events from Z decay and large missing energy linked to dark Higgs decaying to invisible particles.
@@ 2 The simplified Dark Higgs model
The CMS and ATLAS detectors have extensively searched for Z bosons over the years, confirming that heavy neutral gauge bosons do not exist in the mass range of 0.2 to 5.15 TeV. Therefore, we focus on the production of light neutral gauge bosons (Z) below 90 GeV at the FCC-ee.
@@ 3 Simulation of SM backgrounds and signal samples
Table 2: The simulated SM backgrounds generated from electron-positron collisions at the FCC-ee at $\sqrt{s}=240$ GeV. Their corresponding cross-section times branching ratios for each process, and the generation order are presented. Names of these MC samples and the used generators are stated as well.
@@
The DH scenario signal samples and the corresponding SM background samples are privately produced. They have been generated using the WHIZARD event generator 3.1.1 MG5 . The ISR effect was included and interfaced with Pythia 6.24 for the parton shower model and hadronization R34 . For a fast detector simulation of the IDEA detector model idea , the DELPHES package delphes was used. These were generated from electron-positron collisions at the FCC-ee with a 240 GeV center of mass energy, which corresponds to the circumstances of RUN I.
@@ 4.1 Event pre-selection
Figure 2 illustrates the distribution of the dimuon invariant mass for events that passed the pre-selection criteria outlined in table 4, with $M_{\mu^{+}\mu^{-}}<120$ GeV. In this visualization, the red histogram represents the $Z/\gamma$ background, while the cyan histogram indicates the vector boson pair background ($WW$). The green histogram corresponds to the process $ZZ\rightarrow 2\mu 2\nu$, and the yellow histogram depicts the $ZZ\rightarrow 4\mu$ events. These histograms are presented in a stacked format. While the signals of the DH scenario, which have been generated with different masses (from 30 to 80 GeV) of the Z boson and fixing the coupling constants $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$, are represented by different colored lines, and are overlaid.
@@
where $\sqrt{s}$ represents the center-of-mass energy, $E^{\mu^{+}\mu^{-}}$ denotes the energy of the dimuon pair, and $M_{\mu^{+}\mu^{-}}$ refers to the invariant mass of the dimuon pair. To avoid the influence of the Z and SM Higgs boson ($ZH$) events, we conducted our analysis with a recoil mass set below 90 GeV (i.e. $M_{rec}<90$ GeV).
@@ 4.2 Event final selection and efficiencies
The efficiency plots demonstrate that the application of these four stringent selection criteria effectively suppresses the background from the process $Z/\gamma\rightarrow\mu^{+}\mu^{-},\tau\tau$. This approach also minimizes contamination from events such as $WW$, $ZZ(4\mu)$, and $ZZ(2\mu 2\nu)$. Additionally, it ensures that the signal maintains a consistently flat efficiency across high transverse muon momentum ($p^{\mu}_{T}>40$ GeV).
@@ 6 Summary
The Future Circular Collider operating in Electron-Positron collision mode (FCC-ee) is an outstanding facility for uncovering particles that extend beyond the Standard Model (BSM). It offers a distinct signature for pinpointing unknown particles, such as dark matter, extra neutral gauge bosons, and Kaluza-Klein excitations.
Our research investigated the recoil mass distributions with low-mass dimuon pairs within the mono-Z model framework ($M_{rec}<90$ GeV), using simulated Monte Carlo (MC) data samples from the FCC-ee. These MC samples were generated from electron-positron collisions with $\sqrt{s}=240$ GeV, including signal and Standard Model background events. This setup corresponds to what is anticipated for FCC-ee Run 1, which is expected to have an integrated luminosity of 10.8 ab-1.
In this study, we investigated the effects of a simplified model scenario known as the dark Higgs (DH), focusing on dark Higgs production associated with a $Z^{\prime}$ boson at the FCC-ee. We analyzed the results from the muonic decay mode of the $Z^{\prime}$ boson, with coupling constants fixed at $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$.
We implemented effective discrimination cuts that entirely suppressed the $Z/\gamma$ background and significantly reduced the diboson $WW$ and $ZZ$ contaminations. This approach enabled us to differentiate more effectively between signal events and the SM backgrounds. Consequently, we achieved a noteworthy decrease in the SM backgrounds while maintaining the signal strength by applying the appropriate cuts, which are outlined in Table 4 for the DH scenario.
@@
Finally, if the dark Higgs signal is not observed at the FCC-ee, we set upper limits on the mass of dark Higgs ($M_{h_{D}}$) at the 95% confidence level for the charged muonic decay channel of $Z^{\prime}$. Limits have been established for the dark Higgs scenario with $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$, which excludes the dark Higgs mass range from 20 to 80 GeV.
This study highlights the promising opportunities to directly investigate the existence of a dark Higgs particle within the framework of physics that extends beyond the Standard Model at the FCC-ee. This is especially pertinent for the mono-$Z^{\prime}$ model, which features coupling parameters of $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$, along with dark Higgs masses below 90 GeV. Notably, this domain has remained unexplored by the LHC.

### v2

# Search for the production of dark Higgs in the framework of Mono-Z portal at the FCC-$e$ simulated electron-positron collisions at $\sqrt{s}=240$ GeV

@@ Abstract
In the present work, we study the possible production of the dark Higgs boson ($h_{D}$) candidates, which originated from a simplified-model scenario based on the Mono-Z model, in association with a neutral gauge boson (Z). This study has been performed by studying events with dimuon plus missing transverse energy produced in the simulated electron-positron collisions at the foreseen Future Circular Collider in the Electron-Positron collision mode (FCC-$e$), operating at 240 GeV center of mass energy and integrated luminosity of 10.8 ab-1. In case no new physics has been discovered, we set upper limits at a 95% confidence level on the mass of the dark Higgs.
arxiv: arXiv:7228004 [hep-ph]
@@ 1 Introduction
If the Z does not interact with quarks, the HL-LHC and future hadron colliders will be unable to confirm its existence. In such a scenario, electron-positron colliders, like the proposed Future Circular Collider in its Electron-Positron collision mode (FCC-$e$), will play a vital role. The FCC-$e$ is expected to begin operations at a center of mass energy of 240 GeV ($\sqrt{s}$), with plans for a future upgrade to 365 GeV FCC-ee1 ; FCC-ee2 ; FCC-ee3 .
The initial phase of the FCC-$e$ is designed to be a high-luminosity and high-precision electron-positron collider, functioning as a factory for electroweak processes, top quarks, and particularly, the Higgs boson FCC-ee4 . It provides controllable energy levels and minimizes QCD background noise, making it a crucial tool in this research area.
This analysis explores a light dark Higgs ($h_{D}$) which is produced at FCC-$e$ in addition to a light neutral gauge bosons (Z) with a mass $M_{Z^{\prime}}\leq 90$ GeV, based on the dark Higgs (DH) simplified model within the mono-Z framework R1 . We examine simulated electron-positron collisions at the FCC-$e$ with 240 GeV center of mass energy, focusing on dimuon events from Z decay and large missing energy linked to dark Higgs decaying to invisible particles.
@@ 2 The simplified Dark Higgs model
The CMS and ATLAS detectors have extensively searched for Z bosons over the years, confirming that heavy neutral gauge bosons do not exist in the mass range of 0.2 to 5.15 TeV. Therefore, we focus on the production of light neutral gauge bosons (Z) below 90 GeV at the FCC-$e$.
@@ 3 Simulation of SM backgrounds and signal samples
Table 2: The simulated SM backgrounds generated from electron-positron collisions at the FCC-$e$ at $\sqrt{s}=240$ GeV. Their corresponding cross-section times branching ratios for each process, and the generation order are presented. Names of these MC samples and the used generators are stated as well.
@@
The DH scenario signal samples and the corresponding SM background samples are privately produced. They have been generated using the WHIZARD event generator 3.1.1 MG5 . The ISR effect was included and interfaced with Pythia 6.24 for the parton shower model and hadronization R34 . For a fast detector simulation of the IDEA detector model idea , the DELPHES package delphes was used. These were generated from electron-positron collisions at the FCC-$e$ with a 240 GeV center of mass energy, which corresponds to the circumstances of RUN I.
@@ 4.1 Event pre-selection
Figure 2 illustrates the distribution of the dimuon invariant mass for events that passed the pre-selection criteria outlined in table 4, with $M_{\mu^{+}\mu^{-}}<120$ GeV. In this visualization, the red histogram represents the $Z/\gamma$ background, while the cyan histogram indicates the vector boson pair background (WWWW). The green histogram corresponds to the process $ZZ\rightarrow 2\mu 2\nu$, and the yellow histogram depicts the $ZZ\rightarrow 4\mu$ events. These histograms are presented in a stacked format. While the signals of the DH scenario, which have been generated with different masses (from 30 to 80 GeV) of the Z boson and fixing the coupling constants $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$, are represented by different colored lines, and are overlaid.
@@
where $\sqrt{s}$ represents the center-of-mass energy, $E^{\mu^{+}\mu^{-}}$ denotes the energy of the dimuon pair, and $M_{\mu^{+}\mu^{-}}$ refers to the invariant mass of the dimuon pair. To avoid the influence of the Z and SM Higgs boson (ZHZH) events, we conducted our analysis with a recoil mass set below 90 GeV (i.e. $M_{rec}<90$ GeV).
@@ 4.2 Event final selection and efficiencies
The efficiency plots demonstrate that the application of these four stringent selection criteria effectively suppresses the background from the process $Z/\gamma\rightarrow\mu^{+}\mu^{-},\tau\tau$. This approach also minimizes contamination from events such as WWWW, $ZZ(4\mu)$, and $ZZ(2\mu 2\nu)$. Additionally, it ensures that the signal maintains a consistently flat efficiency across high transverse muon momentum ($p^{\mu}_{T}>40$ GeV).
@@ 6 Summary
The Future Circular Collider operating in Electron-Positron collision mode (FCC-$e$) is an outstanding facility for uncovering particles that extend beyond the Standard Model (BSM). It offers a distinct signature for pinpointing unknown particles, such as dark matter, extra neutral gauge bosons, and Kaluza-Klein excitations.
Our research investigated the recoil mass distributions with low-mass dimuon pairs within the mono-Z model framework ($M_{rec}<90$ GeV), using simulated Monte Carlo (MC) data samples from the FCC-$e$. These MC samples were generated from electron-positron collisions with $\sqrt{s}=240$ GeV, including signal and Standard Model background events. This setup corresponds to what is anticipated for FCC-$e$ Run 1, which is expected to have an integrated luminosity of 10.8 ab-1.
In this study, we investigated the effects of a simplified model scenario known as the dark Higgs (DH), focusing on dark Higgs production associated with a $Z^{\prime}$ boson at the FCC-$e$. We analyzed the results from the muonic decay mode of the $Z^{\prime}$ boson, with coupling constants fixed at $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$.
We implemented effective discrimination cuts that entirely suppressed the $Z/\gamma$ background and significantly reduced the diboson WWWW and ZZZZ contaminations. This approach enabled us to differentiate more effectively between signal events and the SM backgrounds. Consequently, we achieved a noteworthy decrease in the SM backgrounds while maintaining the signal strength by applying the appropriate cuts, which are outlined in Table 4 for the DH scenario.
@@
Finally, if the dark Higgs signal is not observed at the FCC-$e$, we set upper limits on the mass of dark Higgs ($M_{h_{D}}$) at the 95% confidence level for the charged muonic decay channel of $Z^{\prime}$. Limits have been established for the dark Higgs scenario with $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$, which excludes the dark Higgs mass range from 20 to 80 GeV.
This study highlights the promising opportunities to directly investigate the existence of a dark Higgs particle within the framework of physics that extends beyond the Standard Model at the FCC-$e$. This is especially pertinent for the mono-$Z^{\prime}$ model, which features coupling parameters of $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$, along with dark Higgs masses below 90 GeV. Notably, this domain has remained unexplored by the LHC.